Calculate the heat capacity of 20kg of brass (specific heat capacity of brass = 380Jkg -

1 K -1 )


1
Expert's answer
2021-10-25T15:24:53-0400

The heat capacity is given as follows:


"C = mc"

where "m = 20kg" is the mass of the block and "c = 380\dfrac{J}{kg\cdot K}" is the specific heat capacity of brass. Thus, obtain:


"C = 20kg\cdot 380\dfrac{J}{kg\cdot K} = 7600\dfrac{J}{K}"

Answer. 7600 J/K.